# Electric and spin current vortices in altermagnets

### v1

@@ Abstract
The results of the lattice calculations in the disordered regime show a good qualitative agreement with the current distribution obtained in the kinetic approach, cf. Figs. 2(c) and 4. While the side vortices are hard to resolve, the central vortices look similar in both approaches. Vortical currents can also be observed in the ballistic transport regime, albeit the corresponding vorticity is not unique to altermagnets [39].
@@
Figure 4: Spin current distribution obtained via the Keldysh formalism presented in the Appendix. The contacts of the width $w=10a$ are at $x=0,L_{x}$ surfaces. We use $60\times 60$ lattice with the following parameters: $t_{m}=0.3t$, $\mu=-2.0t$, and the voltage bias $eV=0.2t$.
@@
Material estimates.— The characteristic values of the key quantities, i.e., magnetic field and current, are
@@
To provide an estimate of the observed effects, we use the material parameters of such altermagnetic candidates as RuO2  [^4] While the recent studies suggest that RuO2 is nonmagnetic [62, 63, 64], it may support altermagnetism in thin films due to strain [65, 66]. and KRu4O8. We use $t_{0}a^{-2}=2.5~{}\mbox{eV}$, $t_{1}/t_{0}=t_{2}/t_{0}=0.2$ and $\mu=0.4~{}\mbox{eV}$ for RuO2 [4, 55, 56] as well as $t_{1}/t_{0}=t_{2}/t_{0}=0.36$ and $\mu=0.05~{}\mbox{eV}$ for KRu4O8 [9, 56].
[…]
Summary and conclusion.— In this Letter, we showed that the altermagnetic spin splitting in finite-size samples with point contacts allows for swirling electric and spin currents even in disordered systems, i.e., in the Ohmic transport regime. The shape of the vortices encodes the relative orientation of the sample boundaries and the altermagnetic nodal planes, see Fig. 2. Swirling electric (spin) currents can be generated if the spin imbalance (voltage) is applied to the contacts, as is supported by Eqs. (7) and (8). The nontrivial distribution of the electric current is manifested in the induced magnetic field and is within reach of modern magnetometry techniques.
Comparing semiclassical calculations in a continuum model with the Keldysh approach on a lattice, we see qualitatively similar results supporting the model-independence of our findings, cf. Figs. 2(c) and 4. We emphasize that our results do not require strong electron-electron interaction (hydrodynamic transport regime) or ultrapure samples (ballistic transport regime), and take place in the Ohmic transport regime with a fully local response to external stimuli. This broadens the selection of material platforms.
@@ Acknowledgements.
Acknowledgments. A.A.H. acknowledges support from the Project No. 0122U000887 of the Department of Physics and Astronomy of the NAS of Ukraine. The work of E.V.G. was supported by the Program “Dynamics of particles and collective excitations in high-energy physics, astrophysics and quantum macrosystems” (No. 0121U109612) of the Department of Physics and Astronomy of the NAS of Ukraine. This work was supported by the Research Council of Norway through Grant No. 353894 and its Centres of Excellence funding scheme Grant No. 262633 “QuSpin.” Support from Sigma2 - the National Infrastructure for High Performance Computing and Data Storage in Norway, project NN9577K, is acknowledged. P.O.S. acknowledges useful communications with Alireza Qaiumzadeh and Xanthe Verbeek, as well as is grateful to the QuSpin center of excellence for warm hospitality.

### v2

@@ A $d_{xy}$-altermagnet is modeled by the following tight-binding Hamiltonian:
The results of the lattice calculations in the disordered regime show a good qualitative agreement with the current distribution obtained in the kinetic approach; cf. Figs. 2(c) and 4. While the side vortices are hard to resolve due to the low magnitude of the current there, the central vortices look similar in both approaches. Vortical currents can also be observed in the ballistic transport regime, albeit the corresponding vorticity is not unique to altermagnets (see the SM [46]).
@@
Figure 4: Spin current distribution obtained via the Keldysh formalism. The contacts of the width w=10aw=10\,a are at $x=0,L_{x}$ surfaces. We use $60\times 60$ lattice with the following parameters: $t_{m}=0.3\,t$, $\mu=-2.0\,t$, and the voltage bias eV=0.2teV=0.2\,t.
@@
Material estimates. The characteristic values of the key quantities, i.e., magnetic field and current, are
@@
To estimate the observed effects, we use the material parameters of altermagnetic candidates such as RuO2  [^4] While the recent studies suggest that RuO2 is nonmagnetic [76, 77, 78], it may support altermagnetism in thin films due to strain [79, 19]. and KRu4O8. We use $t_{0}a^{-2}=2.5~\mbox{eV}$, $t_{1}/t_{0}=t_{2}/t_{0}=0.2$, and $\mu=0.4~\mbox{eV}$ for RuO2 [4, 63, 64] and $t_{1}/t_{0}=t_{2}/t_{0}=0.36$ and $\mu=0.05~\mbox{eV}$ for KRu4O8 [9, 64].
Assuming the width of ribbon $L=5~\mu\mbox{m}$, typical scattering time $\tau=1~\mbox{ps}$, and fixing the applied potential difference $\delta\phi=1~\mbox{mV}$, we estimate $B_{0}\approx 6~\mu\mbox{T}$ and the current density $j_{0}\approx 24\,\mu\mbox{A/m}$ for RuO2 as well as $B_{0}\approx 1~\mu\mbox{T}$ and the current density $j_{0}=3\,\mu\mbox{A/m}$ for KRu4O8. These values are small but are within reach of modern magnetic field imaging techniques such as nanowire magnetic force microscopy [65, 66], the scanning superconducting quantum interference device (SQUID) magnetometry [67, 66], and also at the edge of the sensitivity of the quantum spin magnetometry [68, 38, 66, 69]. In addition to the magnetic field, one can measure the local distribution of the electric potential, albeit the latter is less sensitive to the nontrivial distribution of currents [46].
Summary and conclusion. In this Letter, we showed that the altermagnetic spin splitting in finite-size samples with point contacts allows for swirling electric and spin currents even in disordered systems; i.e., in the Ohmic transport regime. The shape of the vortices encodes the relative orientation of the sample boundaries and the altermagnetic nodal planes, see Fig. 2. Swirling electric (spin) currents can be generated if the spin imbalance (voltage) is applied to the contacts, as is supported by Eqs. (13) and (14). The nontrivial distribution of the electric current is manifested in the induced magnetic field and is within reach of modern magnetometry, complementing previous transport measurements.
Comparing semiclassical calculations in a continuum model with the Keldysh approach on a lattice, we see qualitatively similar results supporting the model-independence of our findings; cf. Figs. 2(c) and 4. We emphasize that our results do not require strong electron-electron interaction (hydrodynamic transport regime) or ultrapure samples (ballistic transport regime), and take place in the Ohmic transport regime with a fully local response to external stimuli. This broadens the selection of material platforms.
@@ Acknowledgements.
Acknowledgments. A.H. acknowledges support from Project No. 0122U000887 of the Department of Physics and Astronomy of the NAS of Ukraine. The work of E.V.G. was supported by the Program “Dynamics of particles and collective excitations in high-energy physics, astrophysics and quantum macrosystems” (Grant No. 0121U109612) of the Department of Physics and Astronomy of the NAS of Ukraine. This work was supported by the Research Council of Norway through Grant No. 353894 and its Centres of Excellence funding scheme Grant No. 262633 “QuSpin.” Support from Sigma2 - The National Infrastructure for High Performance Computing and Data Storage in Norway, Project NN9577K, is acknowledged. P.O.S. acknowledges useful communications with Alireza Qaiumzadeh and Xanthe Verbeek, and is grateful to the QuSpin center of excellence for warm hospitality.